OpenStudy (anonymous):

find the derivative of f(x)=x/sqrtx in 2 different ways

zepdrix (zepdrix):

Hmm ok. So there are several ways to approach this. Have you tried anything yet? :o

OpenStudy (shamil98):

One way is the quotient rule.

OpenStudy (nincompoop):

use the quotient rule and the product rule

OpenStudy (nincompoop):

basically the quotient rule can be derive using the product rule

OpenStudy (nincompoop):

derived*

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt x } = (\frac{ x }{ \sqrt x } ) \frac{ \sqrt x }{ \sqrt x } = \sqrt x\]

OpenStudy (nincompoop):

product rule's general form is f'(x) = u'v + uv' when applying a quotient into the product rule u/v you will end up with u * v^-1 u(x) * [v(x)]^-1

OpenStudy (nincompoop):

what on earth are you doing @jayz657

OpenStudy (anonymous):

it can be simplified into sqrt(x) and you can just take the derivative of sqrt(x) using the power rule

zepdrix (zepdrix):

True it does. I think he was commenting on the steps you showed. They're really strange :) lol

OpenStudy (anonymous):

i multiplied by 1 since sqrt(x)/sqrt(x) = 1

OpenStudy (nincompoop):

you'll end up using the power rule when you derive the product rule

zepdrix (zepdrix):

Oh ok, so I guess there was another step in there, cancelling the x's. I see :o

OpenStudy (nincompoop):

power rule and chain rule in fact

OpenStudy (nincompoop):

are you suggesting that f'(x) = sqrt (x) is the same as f'(x) = x/sqrt(x) ?

zepdrix (zepdrix):

No. Jay was simplifying the problem before taking the derivative. He was asked to show several ways to take the derivative, so that seems like one nice approach.

OpenStudy (anonymous):

yea i rationalized the denominator and it simplified to sqrt(x)

OpenStudy (nincompoop):

you're still stuck with deriving the same problem utilizing 2 different techniques here are the following: by the definition by product rule (needs chain and power rules) by quotient rule

zepdrix (zepdrix):

\[\Large (\sqrt x)'\quad=\quad (x^{1/2})'\]Why would you need product or quotient rule to find that derivative? :\ Maybe I'm misunderstanding what you're saying.. You just use power rule for that one.

OpenStudy (anonymous):

i would prob just use the quotient rule as the 2nd method

OpenStudy (nincompoop):

it is 1/sqrt(x) in this case which is why I said in my earlier post product rule's general form is f'(x) = u'v + uv' when applying a quotient into the product rule u/v you will end up with u * v^-1 u(x) * [v(x)]^-1

zepdrix (zepdrix):

ya, 1/(2sqrt(x)) :o

OpenStudy (shamil98):

I like how the asker hasn't replied at all LOL

OpenStudy (anonymous):

either method works fine

zepdrix (zepdrix):

XD

OpenStudy (nincompoop):

I know they all work fine the purpose of the exercise is to find out one's proficiency at using all derivative techniques